# MUSE-Net: Missingness-aware mUlti-branching Self-attention Encoder for Irregular Longitudinal Electronic Health Records

### v1

@@ 5 Conclusions
6 Acknowledgement
This research work was supported partially by the National Eye Institute of the National Institutes of Health under Award Number R01EY033861 and partially by the National Heart, Lung, And Blood Institute of the National Institutes of Health under Award Number R01HL172292. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. We also acknowledge the Cerner Corporation and Center for Health Systems Innovation (CHSI) at Oklahoma State University for sharing the Health Facts® EHR database to support this research. Additionally, the authors acknowledge the University of Tennessee High-Performance Scientific Computing group for providing computational resources for this work.
7 Data Availability Statement
The data use agreement with Cerner Health Corporation prohibits authors from redistributing the data. Other researchers, who are interested in obtaining the data used in the study, may directly contact Cerner Heath Corporation for data access.
8 Appendix
8.1 Parameter Specifications in Synthetic Data Generation
Tables 5 and 6 provide the detail information of parameter specification in ARMA models and the feature engineering process.
Table 5: Parameter specifications of the three baseline ARMA models
@@ 8.1 Parameter Specifications in Synthetic Data Generation
Table 6: Feature engineering process to generate time series for each variable
Neg. samples (label = 0) Pos. samples (label = 1)
$m=1$ $\text{ARMA}_{1}$ $\text{ARMA}_{11}=\text{ARMA}_{1}$
$m=2$ $\text{ARMA}_{2}$ $\text{ARMA}_{12}=\text{ARMA}_{2}$
$m=3$ $\text{ARMA}_{3}$ $\text{ARMA}_{13}=\text{ARMA}_{3}$
$m=4$ $\text{ARMA}_{4}=0.8\times\text{ARMA}_{1}$ $\text{ARMA}_{14}=1.1\times\text{ARMA}_{11}$
$m=5$ $\text{ARMA}_{5}=\text{ARMA}_{1}+\text{ARMA}_{2}$ $\text{ARMA}_{15}=\text{ARMA}_{11}+\text{ARMA}_{12}$
$m=6$ $\text{ARMA}_{6}=\text{ARMA}_{3}+\text{ARMA}_{4}$ $\text{ARMA}_{16}=\text{ARMA}_{13}+\text{ARMA}_{14}$
$m=7$ $\text{ARMA}_{7}=\text{ARMA}_{1}+\text{ARMA}_{2}+\text{ARMA}_{3}$ $\text{ARMA}_{7}=\text{ARMA}_{11}+\text{ARMA}_{12}+\text{ARMA}_{13}$
$m=8$ $\text{ARMA}_{8}=\text{ARMA}_{1}\times\text{ARMA}_{2}+\text{ARMA}_{3}\times\text{ARMA}_{4}$ $\text{ARMA}_{18}=\text{ARMA}_{11}\times\text{ARMA}_{12}+\text{ARMA}_{13}\times\text{ARMA}_{14}$
$m=9$ $\text{ARMA}_{9}=\text{ARMA}_{3}+\text{ARMA}_{4}+\text{ARMA}_{1}\times\text{ARMA}_{2}$ $\text{ARMA}_{19}=1.1\times\text{ARMA}_{13}+1.1\times\text{ARMA}_{14}+\text{ARMA}_{11}\times\text{ARMA}_{12}$
$m=10$ $\text{ARMA}_{10}=-\text{ARMA}_{5}+\text{ARMA}_{9}$ $\text{ARMA}_{20}=-\text{ARMA}_{15}+\text{ARMA}_{19}$
8.2 Robust Classifier from Multi-branching Network Models
This appendix shows the proof of Theorem I, which is based on Lemma I and Theorem II.
@@ 8.2 Robust Classifier from Multi-branching Network Models
$D_{F}[x_{1}||x_{2}]\equiv F(x_{1})-F(x_{2})-\langle\nabla F(x_{2}),x_{1}-x_{2}\rangle,~{}~{}~{}~{}~{}x_{1},x_{2}\in\mathcal{X}$ (17)
@@
$\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{F}(g(Y_{p})||f(Y_{p};\boldsymbol{\omega}))]=\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{F}(g(Y_{p})||\bar{f}(Y_{p}))]+\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{F}(\bar{f}(Y_{p})||f(Y_{p};\boldsymbol{\omega}))]$ (18)
@@
(Lemma I) According to the definition of $\bar{f}(Y_{p})$, we first have the following results:
00 0 $=$ $\nabla_{z}\mathbb{E}_{\boldsymbol{\omega}}[D_{F}(z||f(Y_{p};\boldsymbol{\omega}))]|_{z=\bar{f}(Y_{p})}$
$=$ $\nabla_{z}\mathbb{E}_{\boldsymbol{\omega}}[F(z)-F(f(Y_{p};\boldsymbol{\omega}))-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),z-f(Y_{p};\boldsymbol{\omega})\rangle]_{z=\bar{f}(Y_{p})}$
$=$ $\left[\nabla F(z)-\nabla_{z}\mathbb{E}_{\boldsymbol{\omega}}[\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),z\rangle]\right]_{z=\bar{f}(Y_{p})}=[\nabla F(z)-\mathbb{E}_{\boldsymbol{\omega}}[\nabla F(f(Y_{p};\boldsymbol{\omega}))]]_{z=\bar{f}(Y_{p})}$
$\Rightarrow$ $\nabla F(\bar{f}(Y_{p}))=\mathbb{E}_{\boldsymbol{\omega}}[\nabla F(f(Y_{p};\boldsymbol{\omega}))]$ (19)
@@
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[F(g(Y_{p}))-F(\bar{f}(Y_{p}))-\langle\nabla F(\bar{f}(Y_{p})),g(Y_{p})-\bar{f}(Y_{p})\rangle]+$
$\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[F(\bar{f}(Y_{p}))-F(f(Y_{p};\boldsymbol{\omega}))-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),\bar{f}(Y_{p})-f(Y_{p};\boldsymbol{\omega})\rangle]$
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[F(g(Y_{p}))-F(f(Y_{p};\boldsymbol{\omega}))-\langle\nabla F(\bar{f}(Y_{p})),g(Y_{p})-\bar{f}(Y_{p})\rangle-\langle\nabla F(f(\boldsymbol{x};\boldsymbol{\omega})),\bar{f}(Y_{p})-f(Y_{p};\boldsymbol{\omega})\rangle]$
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[F(g(Y_{p}))-F(f(Y_{p};\boldsymbol{\omega}))-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),g(Y_{p})-f(Y_{p};\boldsymbol{\omega})\rangle]$
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{F}(g(Y_{p})||f(Y_{p};\boldsymbol{\omega}))]$
@@
$\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{KL}(P(\boldsymbol{x})||\hat{P}(Y_{p},\boldsymbol{w}))]=\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{KL}(P(Y_{p})||\bar{\hat{P}}(Y_{p}))]+\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}(Y_{p})||\hat{P}(Y_{p};\boldsymbol{\omega}))]$ (21)
where $\bar{\hat{P}}(Y_{p})=\arg\min_{P_{*}}\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(P_{*}(Y_{p})||\hat{P}(Y_{p};\boldsymbol{\omega}))]$, and $D_{KL}(P||\hat{P})=\sum_{j=1}^{J}p_{j}\log\frac{p_{j}}{\hat{p}_{j}}$ is the Kullback-Leibler (KL) divergence. The first term in Eq. (21) characterizes the squared bias, and the second term captures the variance of the classifier model.
@@
(Theorem II) The first important note is that minimizing KL divergence is equivalent to minimizing cross-entropy loss in classification tasks. The cross-entropy loss is defined as
$H(P||\hat{P})=-\sum_{j=1}^{J}p_{j}\log\hat{p}_{j}=D_{KL}(P||\hat{P})-\sum_{j}(p_{j}\log p_{j})$ (22)
where $\sum_{j}(p_{j}\log p_{j})$ is often considered as a constant. Hence, minimizing $H(P||\hat{P})$ is equivalent to minimizing $D_{KL}(P||\hat{P})$. Additionally, KL divergence is a special case of the Bregman divergence. Specifically, if we define the function in Definition I as $F(P)=\sum_{i}p_{i}\log p_{i}$, then
$D_{F}(P||\hat{P})$ $=$ $\sum_{i}(p_{i}\log p_{i})-\sum_{i}(\hat{p}_{i}\log\hat{p}_{i})-\sum_{i}\langle\log\hat{p}_{i}+1,p_{i}-\hat{p}_{i}\rangle$ (23)
$=$ $\sum_{i}(p_{i}\log p_{i})-\sum_{i}(p_{i}\log\hat{p}_{i})=\sum_{i}p_{i}\log\frac{p_{i}}{\hat{p}_{i}}=D_{KL}(P||\hat{P})$
As such, we can leverage Lemma I for the Bregman divergence to prove Theorem II. Specifically, Theorem II follows by defining $F(P)=\sum_{i}p_{i}\log p_{i}$, and replacing $g(Y_{p})$ by the true distribution $P(Y_{P})$ and replacing the prediction function $f(Y_{p};\boldsymbol{\omega})$ by $\hat{P}(Y_{p};\boldsymbol{\omega})$.
@@
(Theorem I) According to Theorem II, the variances of the SB and MB classifiers are defined as (we omit subscripts $Y_{p}$ for notation convenience)
$V_{SB}=\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{SB}(\boldsymbol{\omega}))],~{}~{}~{}V_{MB}=\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{MB}||\hat{P}_{MB}(\boldsymbol{\omega}))]$ (24)
where $\hat{P}_{SB}(\boldsymbol{\omega})$ and $\hat{P}_{MB}(\boldsymbol{\omega})$ are the predicted distribution provided by the SB and MB classifiers, respectively. In the MB model, the predicted probability for class $j$ is $\hat{p}_{MB,j}(\boldsymbol{\omega})=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\hat{p}_{j}^{(i)}(\boldsymbol{\omega})$, where $\hat{p}_{j}^{(i)}(\boldsymbol{\omega})$ is the predicted probability for class $j$ by branching output $i$. Given the assumptions in Theorem I, i.e., $\bar{\hat{P}}^{(i)}=\bar{\hat{P}}_{SB}$ and $\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}^{(i)}||\hat{P}^{(i)}(\boldsymbol{\omega}))]=\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{SB}(\boldsymbol{\omega}))]$ ($i\in\{1,\dots,N_{b}\}$) if all the branches and the SB model are sufficiently trained, we have
$V_{SB}$ $=$ $\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{SB}(\boldsymbol{\omega}))]=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}^{(i)}||\hat{P}^{(i)}(\boldsymbol{\omega}))]$ (25)
$=$ $\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbb{E}_{\boldsymbol{\omega}}\left[\sum_{j=1}^{J}\bar{\hat{p}}_{j}^{(i)}\log\frac{\bar{\hat{p}}_{j}^{(i)}}{\hat{p}_{j}^{(i)}(\boldsymbol{\omega})}\right]=\mathbb{E}_{\boldsymbol{\omega}}\left[\sum_{j=1}^{J}\left(\bar{\hat{p}}_{SB,j}\log\bar{\hat{p}}_{SB,j}-\sum_{i=1}^{N_{b}}\bar{\hat{p}}_{SB,j}\frac{\log\hat{p}_{j}^{(i)}(\boldsymbol{\omega})}{N_{b}}\right)\right]$
@@
$=$ $\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{MB}(\boldsymbol{\omega}))]\geq\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{MB}||\hat{P}_{MB}(\boldsymbol{\omega}))]=V_{MB}$
@@
8.3 Algorithm for GPU Acceleration in MGP Imputation
Recent years have seen remarkable advancements in deep learning [35], largely attributed to new software frameworks that allow easy use of GPUs for efficient optimization [56]. In our study, we utilize GPU acceleration for training MUSE-Net, which has become a fundamental practice in modern deep-learning methodologies. However, using GPUs to speed up MGP imputation remains relatively unexplored and requires further investigation. MGP-based imputation can be computationally intensive because it involves operations of large kernel matrices.
MGP-based imputation involves two steps: training (i.e., learning $\boldsymbol{\Theta}$ by gradient-based algorithms) and inference (i.e., calculating the posterior mean for imputation). Three key components contribute to the computational intensity: multiplication of the inverse kernel matrix with observed values $(\Sigma_{p}^{o})^{-1}\boldsymbol{y}_{p}^{o}$, computation of log determinant $\log|\Sigma_{p}^{o}|$ and trace $\text{Tr}\left((\Sigma_{p}^{o})^{-1}\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\right)$. In most existing MGP imputation techniques [43, 57, 42, 31], calculation of the three quantities hinges on Cholesky decomposition of $(\Sigma_{p}^{o})^{-1}$, which has cubic time complexity and thus poses significant computational demands for large datasets. Furthermore, the nonparallelizable nature of Cholesky decomposition makes it not suitable for GPU acceleration. To address this challenge, we adopt the blackbox matrix-matrix multiplication framework [45] to integrate the modified preconditioned conjugate gradient (mPCG) approach with GPU acceleration into the data imputation workflow.

### v2

@@ V Conclusions
VI Appendix
VI-A Parameter Specifications in Synthetic Data Generation
Tables V and VI provide the parameter specification in ARMA models and the feature engineering process.
TABLE V: Parameter specifications of the three baseline ARMA models
@@ VI-A Parameter Specifications in Synthetic Data Generation
TABLE VI: Feature engineering process to generate time series for each variable.
$m$ Neg. samples (label = 0) Pos. samples (label = 1)
1 $\text{ARMA}_{1}$ $\text{ARMA}_{11}=\text{ARMA}_{1}$
2 $\text{ARMA}_{2}$ $\text{ARMA}_{12}=\text{ARMA}_{2}$
3 $\text{ARMA}_{3}$ $\text{ARMA}_{13}=\text{ARMA}_{3}$
4 $\text{ARMA}_{4}=0.8\times\text{ARMA}_{1}$ $\text{ARMA}_{14}=1.1\times\text{ARMA}_{11}$
5 $\text{ARMA}_{5}=\text{ARMA}_{1}+\text{ARMA}_{2}$ $\text{ARMA}_{15}=\text{ARMA}_{11}+\text{ARMA}_{12}$
6 $\text{ARMA}_{6}=\text{ARMA}_{3}+\text{ARMA}_{4}$ $\text{ARMA}_{16}=\text{ARMA}_{13}+\text{ARMA}_{14}$
7 $\text{ARMA}_{7}=\text{ARMA}_{1}+\text{ARMA}_{2}+\text{ARMA}_{3}$ $\text{ARMA}_{7}=\text{ARMA}_{11}+\text{ARMA}_{12}+\text{ARMA}_{13}$
8 $\text{ARMA}_{8}=\text{ARMA}_{1}\times\text{ARMA}_{2}+\text{ARMA}_{3}\times\text{ARMA}_{4}$ $\text{ARMA}_{18}=\text{ARMA}_{11}\times\text{ARMA}_{12}+\text{ARMA}_{13}\times\text{ARMA}_{14}$
9 $\text{ARMA}_{9}=\text{ARMA}_{3}+\text{ARMA}_{4}+\text{ARMA}_{1}\times\text{ARMA}_{2}$ $\text{ARMA}_{19}=1.1\times\text{ARMA}_{13}+1.1\times\text{ARMA}_{14}+\text{ARMA}_{11}\times\text{ARMA}_{12}$
10 $\text{ARMA}_{10}=-\text{ARMA}_{5}+\text{ARMA}_{9}$ $\text{ARMA}_{20}=-\text{ARMA}_{15}+\text{ARMA}_{19}$
VI-B Robust Classifier from Multi-branching Network Models
This appendix shows the proof of Theorem I.
@@ VI-B Robust Classifier from Multi-branching Network Models
$D_{F}[x_{1}||x_{2}]\equiv F(x_{1})-F(x_{2})-\langle\nabla F(x_{2}),x_{1}-x_{2}\rangle,$
$x_{1},x_{2}\in\mathcal{X}$ (19)
@@
$\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{F}(g(Y_{p})||f(Y_{p};\boldsymbol{\omega}))]=$
@@
(Lemma I) From the definition of $\bar{f}(Y_{p})$, we have:
00 0 $=$ $\nabla_{z}\mathbb{E}_{\boldsymbol{\omega}}[D_{F}(z||f(Y_{p};\boldsymbol{\omega}))]\Big{|}_{z=\bar{f}(Y_{p})}$ (21)
$=$ $\nabla_{z}\mathbb{E}_{\boldsymbol{\omega}}\Big{[}F(z)-F(f(Y_{p};\boldsymbol{\omega}))$
$-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),z-f(Y_{p};\boldsymbol{\omega})\rangle\Big{]}\Big{|}_{z=\bar{f}(Y_{p})}$
$=$ $\Big{[}\nabla F(z)-\nabla_{z}\mathbb{E}_{\boldsymbol{\omega}}\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),z\rangle\Big{]}\Big{|}_{z=\bar{f}(Y_{p})}$
$=$ $\Big{[}\nabla F(z)-\mathbb{E}_{\boldsymbol{\omega}}[\nabla F(f(Y_{p};\boldsymbol{\omega}))]\Big{]}\Big{|}_{z=\bar{f}(Y_{p})}$
$\Rightarrow$ $\nabla F(\bar{f}(Y_{p}))=\mathbb{E}_{\boldsymbol{\omega}}[\nabla F(f(Y_{p};\boldsymbol{\omega}))].$
@@
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}\Big{[}F(g(Y_{p}))-F(\bar{f}(Y_{p}))$
$\quad-\langle\nabla F(\bar{f}(Y_{p})),g(Y_{p})-\bar{f}(Y_{p})\rangle\Big{]}$
$+\mathbb{E}_{Y_{p},\boldsymbol{\omega}}\Big{[}F(\bar{f}(Y_{p}))-F(f(Y_{p};\boldsymbol{\omega}))$
$\quad-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),\bar{f}(Y_{p})-f(Y_{p};\boldsymbol{\omega})\rangle\Big{]}$
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}\Big{[}F(g(Y_{p}))-F(f(Y_{p};\boldsymbol{\omega}))$
$\quad-\langle\nabla F(\bar{f}(Y_{p})),g(Y_{p})-\bar{f}(Y_{p})\rangle$
$\quad-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),\bar{f}(Y_{p})-f(Y_{p};\boldsymbol{\omega})\rangle\Big{]}$
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}\Big{[}F(g(Y_{p}))-F(f(Y_{p};\boldsymbol{\omega}))$
$\quad-\langle\nabla F(f(Y_{p};\boldsymbol{\omega})),g(Y_{p})-f(Y_{p};\boldsymbol{\omega})\rangle\Big{]}$
$=$ $\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{F}(g(Y_{p})||f(Y_{p};\boldsymbol{\omega}))].$
@@
$\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{KL}(P(\boldsymbol{x})||\hat{P}(Y_{p},\boldsymbol{w}))]=\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{KL}(P(Y_{p})||\bar{\hat{P}}(Y_{p}))]$
$+\mathbb{E}_{Y_{p},\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}(Y_{p})||\hat{P}(Y_{p};\boldsymbol{\omega}))].$ (23)
where $\bar{\hat{P}}(Y_{p})=\arg\min_{P_{*}}\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(P_{*}(Y_{p})||\hat{P}(Y_{p};\boldsymbol{\omega}))]$ and $D_{KL}(P||\hat{P})=\sum_{j=1}^{J}p_{j}\log\frac{p_{j}}{\hat{p}_{j}}$ is the Kullback-Leibler (KL) divergence. The first term in Eq. (23) is the squared bias, and the second term captures the variance of the classifier model.
@@
(Theorem II) Note that minimizing KL divergence is equivalent to minimizing cross-entropy loss in classification tasks. The cross-entropy loss is defined as $H(P||\hat{P})=-\sum_{j=1}^{J}p_{j}\log\hat{p}_{j}=D_{KL}(P||\hat{P})-\sum_{j}(p_{j}\log p_{j})$, where $\sum_{j}(p_{j}\log p_{j})$ is often considered as a constant. Hence, minimizing $H(P||\hat{P})$ is equivalent to minimizing $D_{KL}(P||\hat{P})$. Additionally, KL divergence is a special case of the Bregman divergence. Specifically, if we define the function in Definition I as $F(P)=\sum_{i}p_{i}\log p_{i}$, then
$D_{F}(P||\hat{P})$ (24)
$=$ $\sum_{i}(p_{i}\log p_{i})-\sum_{i}(\hat{p}_{i}\log\hat{p}_{i})-\sum_{i}\langle\log\hat{p}_{i}+1,p_{i}-\hat{p}_{i}\rangle$
$=$ $\sum_{i}p_{i}\log\frac{p_{i}}{\hat{p}_{i}}=D_{KL}(P||\hat{P})$
As such, we can use Lemma I for the Bregman divergence to prove Theorem II by defining $F(P)=\sum_{i}p_{i}\log p_{i}$, and replacing $g(Y_{p})$ by the true distribution $P(Y_{P})$ and replacing the prediction $f(Y_{p};\boldsymbol{\omega})$ by $\hat{P}(Y_{p};\boldsymbol{\omega})$. ∎
@@
(Theorem I) From Theorem II, the variances of SB and MB classifiers are (omitting subscripts $Y_{p}$ for notation convenience)
$V_{SB}=\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{SB}(\boldsymbol{\omega}))],$ (25)
$V_{MB}=\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{MB}||\hat{P}_{MB}(\boldsymbol{\omega}))]$
where $\hat{P}_{SB}(\boldsymbol{\omega})$ and $\hat{P}_{MB}(\boldsymbol{\omega})$ are the predictions provided by the SB and MB classifiers. In the MB model, the predicted probability for class $j$ is $\hat{p}_{MB,j}(\boldsymbol{\omega})=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\hat{p}_{j}^{(i)}(\boldsymbol{\omega})$, where $\hat{p}_{j}^{(i)}(\boldsymbol{\omega})$ is the predicted probability for class $j$ by branch $i$. Given the assumptions in Theorem I, i.e., $\bar{\hat{P}}^{(i)}=\bar{\hat{P}}_{SB}$ and $\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}^{(i)}||\hat{P}^{(i)}(\boldsymbol{\omega}))]=\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{SB}(\boldsymbol{\omega}))]$ ($i\in\{1,\dots,N_{b}\}$) if all the branches and the SB model are sufficiently trained, we have
$V_{SB}$ $=$ $\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{SB}(\boldsymbol{\omega}))]$ (26)
$=$ $\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}^{(i)}||\hat{P}^{(i)}(\boldsymbol{\omega}))]$
$=$ $\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbb{E}_{\boldsymbol{\omega}}\left[\sum_{j=1}^{J}\bar{\hat{p}}_{j}^{(i)}\log\frac{\bar{\hat{p}}_{j}^{(i)}}{\hat{p}_{j}^{(i)}(\boldsymbol{\omega})}\right]$
$=$ $\mathbb{E}_{\boldsymbol{\omega}}\left[\sum_{j=1}^{J}\left(\bar{\hat{p}}_{SB,j}\log\bar{\hat{p}}_{SB,j}-\sum_{i=1}^{N_{b}}\bar{\hat{p}}_{SB,j}\frac{\log\hat{p}_{j}^{(i)}(\boldsymbol{\omega})}{N_{b}}\right)\right]$
@@
$=$ $\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}||\hat{P}_{MB}(\boldsymbol{\omega}))]\geq\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{MB}||\hat{P}_{MB}(\boldsymbol{\omega}))]$
$=$ $V_{MB}$
@@
VI-C Algorithm for GPU Acceleration in MGP Imputation
In our study, we utilize GPU acceleration for training MUSE-Net, which is a fundamental practice in modern deep learning. However, using GPUs to speed up MGP imputation remains relatively unexplored and requires further investigation. MGP-based imputation can be computationally intensive because it involves operations of large kernel matrices. Three key components contribute to the computational intensity: multiplication of the inverse kernel matrix with observed values $(\Sigma_{p}^{o})^{-1}\boldsymbol{y}_{p}^{o}$, computation of log determinant $\log|\Sigma_{p}^{o}|$ and trace $\text{Tr}\left((\Sigma_{p}^{o})^{-1}\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\right)$. In most existing MGP imputation techniques [44, 58, 43, 31], calculation of the three quantities hinges on Cholesky decomposition of $(\Sigma_{p}^{o})^{-1}$, which has cubic time complexity. Furthermore, the nonparallelizable nature of Cholesky decomposition makes it not suitable for GPU acceleration. To address this challenge, we adopt the blackbox matrix-matrix multiplication framework [46] to integrate the modified preconditioned conjugate gradient (mPCG) approach with GPU acceleration into the data imputation workflow.